# Multi-AUV Cooperative Underwater Multi-Target Tracking Based on Dynamic-Switching-enabled Multi-Agent Reinforcement Learning

### v2

@@ Abstract
With the rapid development of underwater communication, sensing, automation, robot technologies, autonomous underwater vehicle (AUV) swarms are gradually becoming popular and have been widely promoted in ocean exploration and underwater tracking or surveillance, etc. However, the complex underwater environment poses significant challenges for AUV swarm-based accurate tracking for the underwater moving targets. In this paper, we aim at proposing a multi-AUV cooperative underwater multi-target tracking algorithm especially when the real underwater factors are taken into account. We first give normally modelling approach for the underwater sonar-based detection and the ocean current interference on the target tracking process. Then, we regard the AUV swarm as a underwater ad-hoc network and propose a novel Multi-Agent Reinforcement Learning (MARL) architecture towards the AUV swarm based on Software-Defined Networking (SDN). It enhances the flexibility and scalability of the AUV swarm through centralized management and distributed operations. Based on the proposed MARL architecture, we propose the ”dynamic-attention switching” and ”dynamic-resampling switching” mechanisms, to enhance the efficiency and accuracy of AUV swarm cooperation during task execution. Finally, based on a proposed AUV classification method, we propose an efficient cooperative tracking algorithm called ASMA. Evaluation results demonstrate that our proposed tracking algorithm can perform precise underwater multi-target tracking, comparing with many of recent research products in terms of convergence speed and tracking accuracy.
@@
AUV swarm network, multi-target tracking, dynamic-switching attention, dynamic-switching resampling, software-defined networking.
@@ I Introduction
Covering more than 70% of Earth’s surface, the ocean is a vital life-support system with abundant biological and mineral resources [1, 2, 3]. With the advancement of technology and our increasing knowledge of the oceans, the need for efficient and intelligent exploration technologies, particularly in ocean environment surveillance [4, 5], disaster prevention, and rescue operations [6, 7], becomes increasingly pressing. In these years, the rapid development of underwater robots, underwater communication, underwater sensing technologies, has given birth to the Autonomous Underwater Vehicle (AUV). As of today, AUV has become indispensable tools in a wide range of marine applications especially in underwater target tracking [8, 9], due to its ability to autonomously navigate underwater environments.
@@
With the continuous improvement and development of swarm intelligence theory, the control technology of AUV swarms has been greatly propelled and enhanced [10, 11]. The AUV swarms have demonstrated unique advantages in executing tasks in complex marine environments, e.g., the AUV-swarm-based cooperative underwater data collection [12], cooperative underwater target tracking [13]. Through cooperation manner, the AUVs in AUV swarm can cover broader areas, enhancing the efficiency and accuracy of underwater operatiion. Particularly in underwater cooperative multi-target tracking, the AUV swarms have been demonstrated exceptional capabilities in cooperatively tracking various underwater targets such as underwater biological communities [14] or environmental pollutants [15]. This is crucial not only for underwater scientific research such as ecological [16] surveys but also for various categories of civil [17] and military applications [18], such as supporting maritime search, rescue operations and maritime boundary surveillance, etc.
@@
However, effectively managing and cooperating such a complex multi-agent system, especially in dynamic and unpredictable marine environments, has to face many challenges. Particularly in underwater multi-target tracking, these challenges mainly manifest in several aspects: The complexity and dynamism of the marine environment, such as ocean currents interference [19, 20], this demands AUV swarms to possess high adaptability and flexibility to cope with sudden environmental changes and unknown underwater obstacles; Effective communication is crucial for guaranteeing the AUV swarm-based underwater cooperative operations, but the high latency and low bandwidth characteristics of underwater communication especially underwater acoustic-based communication limit the efficiency of the information exchange among the AUVs in AUV swarm and cannot support flexible swarm cooperation [21, 22, 23]. Thus, it requires a scalable network architecture to re-organize and re-define the information exchange architecture for AUV swarm system; Meanwhile, in reality, due to productivity constraints, AUV swarms often comprise various heterogeneous AUVs, making AUV swarm-based cooperative underwater operation more challenging and impeding support for cooperative swarm intelligence [24]. Further, traditional MARL algorithms suffer from long runtimes, slow convergence speeds, and high resource consumption, making MARL challenging to deploy in large-scale AUV swarm.
@@
    We regard the AUV swarm as a AUV swarm network, utilize SDN to optimize AUV swarm network and enhance the flexibility and scalability of the AUV swarm;
@@
    We propose a multi-agent reinforcement learning algorithm incorporating dynamic-switching mechanism (including both ”dynamic-attention switching” and ”dynamic-resampling switching”) to enhance AUV swarm network’s self-learning efficiency and accuracy;
@@
The rest of this paper is organized as follows: Related works on the paper’s theme is surveyed in Sec. II; Preliminary materials are introduced in Sec. III; The proposed FATHOM-Net is detailed in Sec. IV; The proposed cooperative tracking algorithm is presented in Sec. V; Evaluation results are shown in Sec. VI; The paper concludes in Sec. VII with a discussion of some future research directions derived from this paper.
@@ II Related Works
In this section, we investigate recent advances related to the main research topics, specifically categorized into the following three areas: 1) SDN-based underwater multi-agent system, 2) Single AUV-based target tracking, and 3) AUV swarm-based target tracking.
@@ II-C AUV swarm-based target tracking
Totally, multi-AUV-based underwater target tracking technology has received widespread research attention. However, this research field still faces several challenges, such as adaptability to dynamic and complex marine environments, flexibility and robustness of AUV swarm networks, and how to stably and accurately track multiple targets. In response to these issues, our work employ the SDN technology to improve the scalability and functionability of the AUV swarm system. On account of the above system architecture, we propose a dynamic-switching-enabled MARL architecture which is more feasible for AUV swarm system’s self-learning. And, we propose as well a novel AUV classification method to improve the scalability of the AUV swarm-based tracking system. In the following, we will prove our work demonstrates precise multi-target tracking in dynamic and complex marine environments, compared to many existing studies.
@@ III-B Ocean current modeling
However, the Navier-Stokes equations involve a large number of partial differential calculations, resulting in high computational costs and long processing times. To reduce the complexity of the model and obtain preliminary fluid dynamic characteristics, we assume that fluid flow is laminar and relatively uniform. Neglecting the nonlinear convective term $\mathbf{u}\cdot\nabla\mathbf{u}$ in the Navier-Stokes equations simplifies it to a linear problem that can be expressed by the following.
$\rho\frac{\partial\mathbf{u}}{\partial t}=-\nabla p+\mu\nabla^{2}\mathbf{u}+\mathbf{F}$ (3)
After calculating $\mathbf{u}$ and $\frac{\partial\mathbf{u}}{\partial t}$ through the Navier-Stokes equations, we use the drag equation (Eq. 4), lift equation (Eq. 5), and virtual mass equation (Eq. 6) to calculate the forces acting on the structure under the influence of ocean currents:
@@
Algorithm 1 Simplified fluid dynamics calculation algorithm
@@
7:         Calculate the velocity time derivative $\frac{\partial\mathbf{u}}{\partial t}$ by simplified Navier-Stokes equation.
@@ IV Hierarchical Software-Defined Multi-AUV Reinforcement Learning Architecture
In this section, we introduce Flexible Autonomous Training Hierarchy for Operational Marine Network (FATHOM-Net), which is a software-defined MARL architecture based on Centralized Training and Decentralized Execution.
Figure 2: Proposed FATHOM-Net
IV-A Overview of FATHOM-Net
FATHOM-Net framework consists of three functional layers: global-control training layer, local-control training layer, and application execution tracking layer, which can be detailed as the following.
@@ IV-A Overview of FATHOM-Net
Overall, FATHOM-Net reduces the coupling between individual AUVs, addressing the scalability issues present in traditional architectures. The FATHOM-Net framework is summarized in Fig. 2. The entire task classification processing is summarized in Fig. 3.
IV-B Beacon framework for FATHOM-Net
In order to ensure precise target tracking based on the proposed FATHOM-Net, it is crucial to establish both local and global network views for LC-AUVs and USV-GC. Therefore, we propose an SDN beacon framework to synchronize and share information among ET-AUVs, LC-AUVs, and USV-GC. The proposed SDN beacon framework can be carried out by the following:
@@ V Proposed Cooperative Tracking Algorithm
In this section, we showcase our proposed cooperative tracking algorithm. Especially, we firstly propose a classification algorithm called AUV Scalable Management and Allocation (ASMA) to classify/allocate ET-AUVs to each formation for performing cooperative multi-target tracking. To enhance the tracking accuracy and the convergence speed of the AUV swarm network, we employ the concept of ”Dynamic-Switching”, integrate the ”Dynamic-Switching Attention” mechanism into the Actor-Critic model and develop the Dymiac-Switching-Based MARL (DSBM) architecture. Additionally, to improve the efficiency of better sample utilization in DSBM, we introduce a novel ”Dynamic-Switching Resampling” mechanism based on experience replay buffer mechanism.
@@ V-A AUV classification method
Figure 3: Task classification based on FATHOM-Net
@@
The rule-based expert system takes the fuzzy membership degrees as input. It then calculates the overall score $\mathit{S}_{i,j}$ acquired by agent $i$ and target $j$. The notations and their meanings are shown in Table I.
@@
V-B Proposed dynamic-switching-based MARL
Normally, in MARL, there are two major challenges: the high variance problem of traditional strategy estimation methods and the slow convergence problem caused by the high dimensionality of Critic networks in multi-agent systems. To address these issues, in this paper, we introduce Temporal Difference techniques to reduce evaluation variance and propose DSBM, which adopts the ”Dynamic-Switching” mechanism to improve learning efficiency and accelerate policy convergence. The ”Dynamic-Switching” mechanism focuses on optimizing the handling of high-value samples, as shown on the right side of Fig. 2.
$AUV_{\text{be}}=\text{argmax}(R_{f})$ (15)
where $R_{f}$ is the set of reward value for the AUVs at round $f$.
The specific selection process is as follows:
  1. 1.
    If the current ET-AUV has the highest reward, then randomly select two from the remaining ET-AUVs to avoid the algorithm getting stuck in local optima.
  2. 2.
$AUV_{\text{ord}}=R_{f}\backslash AUV_{\text{ad}}$ (17)
$m_{1}=cat_{k\in AUV_{ad}}(o_{k},a_{k})$ (18)
$m_{2}=\sum_{k\in AUV_{ord}}\omega_{k}f(o_{k},a_{k})$ (19)
$\omega_{k}=\frac{r_{fk}}{\sum_{m\in AUV_{ord}}r_{fm}}$ (20)
where $r_{fk}$ is the reward of $k$-th ET-AUV at $f$-th round.
$Q^{\psi}_{i}(o,a)=f_{i}(m_{1},m_{2})$ (21)
where $f_{i}$ is a three-layer MLP.
Our approach involves extracting a subset $D_{1}$ comprising $n$ samples from $D$ in each training round, which is then used to update the Critic network. Subsequently, the update process selects the top 50% highest-reward samples from $D_{1}$. These selected samples form a new set $D_{2}$, whose samples are then used for a second update for the Critic network. This mechanism ensures that the focus of network learning is on the most valuable experiences.
$\hat{\omega}\leftarrow\tau\omega+(1-\tau)\hat{\omega}$ (22)
where
To slightly adjust the policy of each agent, we utilize gradient ascent to update their policy parameters (Eq. 25).
where $a_{i}=\pi_{\theta_{i}}(o_{i})$.
V-C Proposed cooperative tracking policy
In this section, we introduce the proposed reward function. In reinforcement learning, by modeling the problem as MDP, the problem can be transformed into a corresponding maximization of the cumulative expected reward. Therefore, we define the reward function as follows (Eq. 26):
@@ V-C Proposed cooperative tracking policy
For ET-AUV $i$, the tracking reward $r_{i}^{\delta}$ is designed to control the distance between EC-AUV and the tracking target, thereby maintaining the agent and the target within an best tracking distance range (Eq. 27).
@@
where $v_{\text{age}}(i)$ is the velocity of the agent, $v_{\text{tar}}(j)$ is the velocity of the target.

### v3

@@ Abstract
In recent years, autonomous underwater vehicle (AUV) swarms are gradually becoming popular and have been widely promoted in ocean exploration or underwater tracking, etc. In this paper, we propose a multi-AUV cooperative underwater multi-target tracking algorithm especially when the real underwater factors are taken into account. We first give normally modelling approach for the underwater sonar-based detection and the ocean current interference on the target tracking process. Then, based on software-defined networking (SDN), we regard the AUV swarm as a underwater ad-hoc network and propose a hierarchical software-defined multi-AUV reinforcement learning (HSARL) architecture. Based on the proposed HSARL architecture, we propose the ”Dynamic-Switching” mechanism, it includes ”Dynamic-Switching Attention” and ”Dynamic-Switching Resampling” mechanisms which accelerate the HSARL algorithm’s convergence speed and effectively prevents it from getting stuck in a local optimum state. Additionally, we introduce the reward reshaping mechanism for further accelerating the convergence speed of the proposed HSARL algorithm in early phase. Finally, based on a proposed AUV classification method, we propose a cooperative tracking algorithm called Dynamic-Switching-Based MARL (DSBM)-driven tracking algorithm. Evaluation results demonstrate that our proposed DSBM tracking algorithm can perform precise underwater multi-target tracking, comparing with many of recent research products in terms of various important metrics.
@@
AUV swarm, multi-target tracking, Dynamic-Switching Attention, Dynamic-Switching Resampling, software-defined networking.
@@ I Introduction
Covering more than 70% of Earth’s surface, the ocean is a vital life-support system with abundant biological and mineral resources [1, 2]. With the advancement of technology and our increasing knowledge of the oceans, the need for efficient and intelligent exploration technologies, particularly in ocean environment surveillance [3, 4], disaster prevention, and rescue operations [5, 6], becomes increasingly pressing. In these years, the rapid development of underwater robots, underwater communication, underwater sensing technologies, has given birth to the Autonomous Underwater Vehicle (AUV). As of today, AUV has become indispensable tools in a wide range of marine applications especially in underwater target tracking [7], due to its ability to autonomously navigate underwater environments.
@@
With the continuous improvement and development of swarm intelligence theory, the control technology of AUV swarms has been greatly propelled and enhanced [8, 9]. The AUV swarms have demonstrated unique advantages in executing tasks in complex marine environments, e.g., the AUV-swarm-based cooperative underwater data collection [10], cooperative underwater target tracking [11]. Through cooperation manner, the AUVs in AUV swarm can cover broader areas, enhancing the efficiency and accuracy of underwater operation. Particularly in underwater cooperative multi-target tracking, the AUV swarms have been demonstrated exceptional capabilities in cooperatively tracking various underwater targets such as underwater biological communities [12] or environmental pollutants [13]. This is crucial not only for underwater scientific research such as ecological [14] surveys but also for various categories of civil [15] and military applications [16], such as supporting maritime search, rescue operations and maritime boundary surveillance, etc. Notably, recently, the multi-agent intelligence or Multi-Agent Reinforcement Learning (MARL) has been emergent and provides an efficient platform for scheduling the AUV swarm to perform cooperative multi-target tracking.
@@
However, effectively managing and cooperating such a complex multi-agent system, especially in dynamic and unpredictable marine environments, has to face many challenges. Particularly in underwater multi-target tracking, these challenges mainly manifest in several aspects: 1) complex marine environment: the complexity and dynamism of the marine environment, such as ocean currents interference [17, 18], this demands AUV swarms to possess high adaptability and flexibility to cope with sudden environmental changes and unknown underwater obstacles; 2) limited underwater communication: the high latency and low bandwidth characteristics of underwater communication, especially underwater acoustic-based communication, limit the efficiency of information exchange among the AUVs in the swarm. These limitations prevent the support of flexible swarm cooperation [19, 20, 21]. Thus, it requires a scalable network architecture to re-organize and re-define the information exchange architecture for AUV swarm system; 3) heterogeneous AUVs: due to productivity constraints, AUV swarms often comprise various heterogeneous AUVs, making AUV swarm-based cooperative underwater operation more challenging and impeding support for cooperative swarm intelligence [22]; 4) scalable deployment: traditional MARL algorithms suffer from long runtime, slow convergence speeds, and high resource consumption, making MARL challenging to deploy in large-scale AUV swarm.
Additionally, in academia, there are two major challenges in MARL: the high variance problem of traditional strategy estimation methods and the slow convergence problem caused by the high dimensionality of Critic networks in multi-agent systems.
Therefore, how to optimize these algorithms to address these challenges without sacrificing efficiency, has become a focal point of current research.
@@
    We regard the AUV swarm as a AUV swarm network, utilize SDN to optimize AUV swarm network and propose a hierarchical software-defined multi-AUV reinforcement learning architecture;
@@
    Based on the proposed MARL architecture, we incorporate Dynamic-Switching mechanism (including both ”Dynamic-Switching Attention” and ”Dynamic-Switching Resampling”) to enhance AUV swarm network’s self-learning efficiency and accuracy;
@@
The rest of this paper is organized as follows. The related works on the paper’s theme is surveyed in Sec. II; In Sec. III, the preliminary materials are showcased; The proposed HSARL is detailed in Sec. IV; The proposed cooperative tracking algorithm is presented in Sec. V; In Sec. VI, various categories of evaluation results are presented; Sec. VII concludes the paper and discusses some future research directions derived from this paper.
@@ II Related Works
In this section, we investigate recent advances related to the main research topics, specifically categorized into the following three areas: 1) SDN-based underwater multi-agent system; 2) Single AUV-based target tracking; 3) AUV swarm-based target tracking; 4) AI/ML in underwater multi-agent distributed strategies.
@@ II-C AUV swarm-based target tracking
II-D AI/ML in underwater multi-agent distributed strategies
Normally, a distributed multi-agent strategy for AUV-based underwater systems involves the coordination of multiple Autonomous Underwater Vehicles (AUVs) to accomplish complex tasks. This approach leverages distributed computing and communication networks, enabling each AUV to share environmental information and decisions in real-time, thereby enhancing the efficiency and accuracy of task execution.
In [34], the authors discuss various AI/ML techniques and their specific applications in 6G, including Supervised Learning, Unsupervised Learning, Deep Learning (DL), and Reinforcement Learning (RL). These methods also find applications in distributed multi-agent strategies for underwater systems, especially Unsupervised Learning and RL.
Unsupervised Learning: In [35], the authors employ the K-Means clustering algorithm for network partitioning and swarm head selection based on network topology and residual energy of nodes. This aims to achieve efficient data aggregation in underwater sensor networks, thereby improving data transmission efficiency and network energy utilization.
Reinforcement Learning (RL): In [10], the authors propose a target uncertainty map assisted data collection scheme for AUV swarms based on the multi-agent proximal policy optimization (MAPPO) algorithm. This approach utilizes current and past search and collection results to establish a target uncertainty map. Combined with artificial potential fields, it ultimately significantly enhances data collection efficiency and accuracy.
Totally, multi-AUV-based underwater target tracking technology has received widespread research attention. However, this research field still faces several challenges, such as adaptability to dynamic and complex marine environments, flexibility and robustness of AUV swarm networks, and how to stably and accurately track multiple targets. In response to these issues, our work employ the SDN technology to improve the scalability and functionality of the AUV swarm system. On account of the above system architecture, we propose a Dynamic-Switching-enabled MARL architecture which is more feasible for AUV swarm system’s self-learning. And, we propose as well a novel AUV classification method to improve the scalability of the AUV swarm-based tracking system.
@@ III-B Ocean current modeling
However, the Navier-Stokes equations involve a large number of partial differential calculations, resulting in high computational costs and long processing times. To reduce the complexity and better handle ocean current, we use the Reynolds-Averaged Navier-Stokes (RANS) equations. The RANS equations are derived from the Navier-Stokes equations by applying a time or statistical averaging process. This averaging effectively reduces computational complexity and has gained widespread use in industrial applications. The RANS equations decompose the velocity into mean and fluctuating components: $u=\overline{u}+u^{\prime}$, where $\overline{u}$ is the mean velocity and $u^{\prime}$ is the fluctuating velocity. The RANS equations can be expressed as follows:
$\rho\left(\frac{\partial\overline{u}}{\partial t}+\overline{u}\cdot\nabla\overline{u}\right)=-\nabla\overline{p}+\mu\nabla^{2}\overline{u}+\nabla\cdot\left(-\rho\overline{u^{\prime}u^{\prime}}\right)+F$ (3)
After calculating $\overline{u}$ and $\frac{\partial\overline{u}}{\partial t}$ through the RANS equations, we use the drag equation (Eq. 4), lift equation (Eq. 5), and virtual mass equation (Eq. 6) to calculate the forces acting on the structure under the influence of ocean currents:
@@
Algorithm 1 RANS-based ocean current interference calculation algorithm
@@
7:         Calculate the velocity time derivative $\frac{\partial\mathbf{u}}{\partial t}$ by RANS equation.
@@ IV Hierarchical Software-Defined Multi-AUV Reinforcement Learning Architecture
In this section, we introduce Hierarchical Software-defined multi-AUV Reinforcement Learning (HSARL), which is a software-defined MARL architecture based on centralized training and decentralized execution.
Figure 2: Proposed HSARL
IV-A Overview of HSARL
HSARL framework consists of three functional layers: global-control training layer, local-control training layer, and application execution tracking layer, which can be detailed as the following.
@@ IV-A Overview of HSARL
Overall, HSARL reduces the coupling between individual AUVs, addressing the scalability issues present in traditional architectures. The HSARL framework is summarized in Fig. 2. The entire task classification processing is summarized in Fig. 3.
IV-B Beacon framework for HSARL
In order to ensure precise target tracking based on the proposed HSARL, it is crucial to establish both local and global network views for LC-AUVs and USV-GC. Therefore, we propose an SDN beacon framework to synchronize and share information among ET-AUVs, LC-AUVs, and USV-GC. The proposed SDN beacon framework can be carried out by the following:
@@ V Proposed Cooperative Tracking Algorithm
In this section, we showcase our proposed cooperative tracking algorithm. Especially, we firstly propose a classification algorithm called AUV Scalable Management and Allocation (ASMA) to classify/allocate ET-AUVs to each formation for performing cooperative multi-target tracking. To enhance the tracking accuracy and the convergence speed of the AUV swarm network, on account of HSARL, we employ the concept of ”Dynamic-Switching”, integrate the ”Dynamic-Switching Attention” mechanism into the Actor-Critic model and develop the Dymiac-Switching-Based MARL (DSBM)-driven tracking algorithm. Additionally, to improve the efficiency of sample utilization in DSBM, we introduce a novel ”Dynamic-Switching Resampling” mechanism based on experience replay buffer mechanism.
@@ V-A AUV classification method
Figure 3: Task classification based on HSARL
@@
The rule-based expert system takes the fuzzy membership degrees as input. It then calculates the overall score $\mathit{S}_{i,j}$ acquired by ET-AUV $i$ and target $j$. The notations and their meanings are shown in Table I.
@@
V-B Proposed cooperative tracking policy
In this section, we introduce the proposed reward function under the HSARL reinforcement learning architecture. In reinforcement learning, by modeling the problem as MDP, the problem can be transformed into a corresponding maximization of the cumulative expected reward. Therefore, we define the reward function as follows (Eq. 15):
@@ V-B Proposed cooperative tracking policy
For ET-AUV $i$, the tracking reward $r_{i}^{\delta}$ is designed to control the distance between EC-AUV and the tracking target, thereby maintaining the ET-AUV and the target within the best tracking distance range (Eq. 16).
@@
where $v_{\text{age}}(i)$ is the velocity of the ET-AUV, $v_{\text{tar}}(j)$ is the velocity of the target.
V-C Proposed DSBM-driven tracking algorithm
V-C1 Proposed DSBM
Normally, as it is aforementioned in Sec. I, there are two major challenges: the high variance problem of traditional strategy estimation methods and the slow convergence problem caused by the high dimensionality of Critic networks in multi-agent systems. To address these issues, in this paper, we introduce temporal difference techniques to reduce evaluation variance and propose DSBM, which adopts the ”Dynamic-Switching” mechanism to improve learning efficiency and accelerate policy convergence. The ”Dynamic-Switching” mechanism focuses on optimizing the handling of high-value samples, as shown on the right side of Fig. 2.
$AUV_{\text{be}}=\text{argmax}(R_{f})$ (20)
where $R_{f}$ is the set of reward value for the AUVs at round $f$.
The specific selection process is as follows:
  1. 1.
    If the current ET-AUV has the highest reward, then randomly select two from the remaining ET-AUVs to avoid the algorithm getting stuck in local optimum state.
  2. 2.
By the following Proof A, it can be seen that the aforementioned part of the proposed Dynamic-Switching Attention mechanism is able to lead to better training results by leveraging high-reward samples for more effective policy updates.
$AUV_{\text{ord}}=R_{f}\backslash AUV_{\text{ad}}$ (22)
$m_{1}=cat_{k\in AUV_{ad}}(o_{k},a_{k})$ (23)
$m_{2}=\sum_{k\in AUV_{ord}}\omega_{k}f(o_{k},a_{k})$ (24)
Proof A To demonstrate that the dynamic-switching attention mechanism accelerates convergence speed, we consider the standard Q-value update for ET-AUV $i$: $Q_{t+1}^{i}(s,a)=Q_{t}^{i}(s,a)+\alpha\delta_{t}^{i}$ (25) where the temporal difference (TD) error is defined as: $\delta_{t}^{i}=r_{t}^{i}+\gamma\max_{a^{\prime}}Q_{t}^{i}(s^{\prime},a^{\prime})-Q_{t}^{i}(s,a)$ (26) With the attention mechanism, the expected TD error becomes a weighted sum of the errors from the best ET-AUV and the others: $\mathbb{E}[\delta_{t}^{i}]=p_{\text{be}}\mathbb{E}[\delta^{\text{be}}]+(1-p_{\text{be}})\mathbb{E}[\delta^{\text{others}}]$ (27) where $p_{\text{be}}$ is the probability of selecting experiences from the best ET-AUV. Since the best ET-AUV has superior performance: $\mathbb{E}[\delta^{\text{be}}]<\mathbb{E}[\delta^{\text{others}}]$ (28) Therefore, increasing $p_{\text{be}}$ leads to a decrease in the expected TD error: $p_{\text{be}}\uparrow\implies\mathbb{E}[\delta_{t}^{i}]\downarrow$ (29) A lower expected TD error results in faster convergence of the Q-values: $\text{Convergence speed}\propto\frac{1}{\mathbb{E}[\delta_{t}^{i}]}$ (30) Thus, the dynamic-switching attention mechanism accelerates the model’s convergence by increasing $p_{\text{be}}$ and reducing $\mathbb{E}[\delta_{t}^{i}]$.
$\omega_{k}=\frac{r_{fk}}{\sum_{m\in AUV_{ord}}r_{fm}}$ (31)
where $r_{fk}$ is the reward of $k$-th ET-AUV at $f$-th round.
$Q^{\psi}_{i}(o,a)=f_{i}(m_{1},m_{2})$ (32)
where $f_{i}$ is a three-layer MLP.
Totally, by the following Proof B, it can be concluded that the proposed dynamic weighting phase in Eq. 31 is able to reduce the variance in policy estimation by dynamically assigning higher weights to high-reward experiences.
Our approach involves extracting a subset $D_{1}$ comprising $n$ samples from $D$ in each training round, which is then used to update the Critic network. Subsequently, the update process selects the top 50% highest-reward samples from $D_{1}$. These selected samples form a new set $D_{2}$, whose samples are then used for a second update for the Critic network.
Proof B Let $Q(s,a)$ be the state-action value function, and $Q_{\text{target}}(s,a)$ be the target value for the update. In traditional Q-learning, the variance of the policy estimation is given by: $\sigma^{2}=\mathbb{E}\left[\left(Q_{\text{target}}(s,a)-\mathbb{E}[Q_{\text{target}}(s,a)]\right)^{2}\right]$ (33) In the Dynamic-Switching Attention mechanism, dynamic weights $\omega_{i}$ are introduced, and the update rule becomes: $Q(s,a)\leftarrow Q(s,a)+\alpha\left(\sum_{i=1}^{N}\omega_{i}Q_{\text{target},i}(s,a)-Q(s,a)\right)$ (34) where $\omega_{i}=\frac{R_{i}}{\sum_{j=1}^{N}R_{j}}$, dynamically adjusted based on the rewards $R_{i}$. The new variance $\sigma^{2}_{\text{new}}$ is: $\sigma^{2}_{\text{new}}=\sum_{i=1}^{N}\omega_{i}^{2}\sigma^{2}_{i}$ (35) Since $\omega_{i}$ is larger for higher rewards $R_{i}$, the contribution of low-value samples is reduced. This leads to: $\sigma^{2}_{\text{new}}<\sigma^{2}_{\text{original}}$ (36) Thus, the dynamic weighting reduces the variance in policy estimation.
$\hat{\omega}\leftarrow\tau\omega+(1-\tau)\hat{\omega}$ (37)
where
To slightly adjust the policy of each ET-AUV, we utilize gradient ascent to update their policy parameters (Eq. 40).
where $a_{i}=\pi_{\theta_{i}}(o_{i})$.
Reward reshaping mechanism: In actual situations, the tracking reward (Eq. 16) plays the most significant role in the reward function. The shorter the distance between the AUV and the target, the smaller the negative reward.
However, in the early tracking stage, the distance between the AUV and the target is too large. As a result, AUV’s any action has a little effect on the system reward, making it easily falls into a local optimum state. Therefore, we introduce a reward reshaping mechanism and propose the reward reshaping term $r_{i}^{\phi}$ (Eq. 41):
$\phi=\frac{\left(p_{\mathrm{tar}}-p_{\mathrm{age}}\right)\cdot v_{\mathrm{age}}}{|p_{\mathrm{tar}}-p_{\mathrm{age}}|\cdot|v_{\mathrm{age}}|}$ (41)
where $p_{\text{tar}}$ is the position vector of the target, $p_{\text{age}}$ is the position vector of the AUV, and $v_{\text{age}}$ is the velocity vector of the AUV.
The reward reshaping term is determined by the angle between the AUV’s velocity direction and the relative direction to the target. The smaller the angle, the higher the reward.
Finally, we introduce the reward reshaping mechanism into the reward function. We obtain the reshaped reward function $Re(s,a)$ as follows:
$Re(s,a)=R_{i}(s,a)+\phi$ (42)
V-C2 Proposed cooperative tracking algorithm